OpenStudy (anonymous):

how would I factor f(x)=x^3-x^2-3x+3 into a linear factor?

OpenStudy (cwrw238):

try grouping: rearranging f(x) = x^3 - 3x - x^2 + 3 = x(x^2 - 3) - 1(x^2 - 3) can you continue?

OpenStudy (anonymous):

well the way I have in my notes has me find this:

OpenStudy (anonymous):